OpenStudy (anonymous):

{I'll give you a medal and become a fan <3 :) please helppp } 8. The ordered pairs (1, 2), (2, 4), (3, 8), (4, 16), and (5,32) represent a function. What is a rule that represents this function? a. y= x^2 b. y= 2x c.y= 2^x d. y= x+2

TheSmartOne (thesmartone):

it is C.

TheSmartOne (thesmartone):

Because y=2^(1)=2 Thus the point (1,2) y=2^2=4 thus (2,4) y=2^3=8 thus (3,8) y=2^4=16 thus (4,16) y=2^5=32 thus (5,32)

TheSmartOne (thesmartone):

@Mimi_Styles

OpenStudy (anonymous):

:O omg! thank you so much!

TheSmartOne (thesmartone):

Your welcome

OpenStudy (anonymous):

may I ask another question, if that's fine? :3

TheSmartOne (thesmartone):

Sure :)

OpenStudy (anonymous):

yay! :D

OpenStudy (anonymous):

okay, give me a minute so I can write it :)

TheSmartOne (thesmartone):

Sure.

OpenStudy (anonymous):

here. I just screenshot it :)

TheSmartOne (thesmartone):

You need to plug in all the ordered pairs into the equations and see what equation works for all the ordered pair.

TheSmartOne (thesmartone):

In this case it would be A.

OpenStudy (anonymous):

ohhh. i just tried in my calculator so I could get it. :) thank youuuu again <3

TheSmartOne (thesmartone):

YW

OpenStudy (anonymous):

I'm sorry if I'm bothering you. am I bother you ?

TheSmartOne (thesmartone):

No your not...

OpenStudy (anonymous):

oki. just checking. ^_^ I didn't wanna bother you if you are busy or anything

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

may I ask another one? :o

TheSmartOne (thesmartone):

Sure.. :)

OpenStudy (anonymous):

yeppie :D oki

TheSmartOne (thesmartone):

Its c because .4 per pound so .4n and the .25 is constant...

TheSmartOne (thesmartone):

@Mimi_Styles

OpenStudy (anonymous):

oooooo oki. ^_^